Question 1 Your answer is partially correct. Try again. Grouper Industries is considering the purchase of new equipment costiplease show work!
Answer #1

Annual profit = (12-7)*17400-52200 = $34800

Net present value = present value of annual profit +present value of residual value - Net cash outflows

= 34800*1.808 + 116000*0.873 - (696000-104400)

= - 427362